An angle measures 36° more than the measure of a supplementary angle. What is the measure of each angle ?

Answer by fractalier(6550)

If we call the angle x, its supplement must be 180-x. Then we write
x = (180-x) + 36
Now solve this...we get
x = 180 - x + 36
2x = 216
x = 108
180-x = 72
The angles are 108 and 72 degrees.